OpenStudy (anonymous):

y = x2 + 4x + 7 How do i write this quadratic function in vertex form?

ganeshie8 (ganeshie8):

rewrite the second term as 2(2x)

ganeshie8 (ganeshie8):

\(y = x^2 + 4x + 7\) \(y = x^2 + 2(2)x + 7\)

ganeshie8 (ganeshie8):

remember the formula for (a + b)^2 ?

OpenStudy (anonymous):

oh, got it! thanks

ganeshie8 (ganeshie8):

good :) so whats the vertex form ha ?

OpenStudy (anonymous):

oh i guess i don't never mind haha